### v1

# On the Convergence Rate of LoRA Gradient Descent

Abstract
The low-rank adaptation (LoRA) algorithm for fine-tuning large models has grown popular in recent years due to its remarkable performance and low computational requirements. LoRA trains two “adapter” matrices that form a low-rank representation of the model parameters, thereby massively reducing the number of parameters that need to be updated at every step. Although LoRA is simple, its convergence is poorly understood due to the lack of Lipschitz smoothness, a key condition for classic convergence analyses. As a result, current theoretical results only consider asymptotic behavior or assume strong boundedness conditions which artificially enforce Lipschitz smoothness. In this work, we provide for the first time a non-asymptotic convergence analysis of the original LoRA gradient descent algorithm, which reflects widespread practice, without such assumptions. Our work relies on three key steps: i) reformulating the problem in terms of the outer product of the stacked adapter matrices, ii) a modified descent lemma for the “Lipschitz-like” reparametrized function, and iii) controlling the step size. With this approach, we prove that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, where $T$ is the number of iterations.
1 Introduction
Modern applications of large language models (LLMs) typically involve self-supervised pretraining of a large foundation model, which is later fine-tuned on a smaller task-specific dataset through supervised learning. Parameter efficient fine-tuning methods, which only train a small number of additional parameters, can efficiently adapt LLMs to a multitude of downstream tasks while maintaining comparable performance to fine-tuning of the full model parameters. One highly popular method is LoRA, or Low-Rank Adaptation [hu2022lora], which trains a low-rank representation of the model parameters. Specifically, once the pretraining stage yields a model weight matrix $W_{0}$, LoRA only updates low-rank adapter matrices $A$, $B$, such that the final weights are $W_{0}+BA$. The matrices $A$ and $B$ can be updated with any optimization method, including gradient descent. However, even this simple algorithm is challenging to analyze; as observed in [sun2024improving] and [malinovsky2024randomizedasymmetricchainlora], even if the original loss function is Lipschitz smooth, the new loss function— reparametrized in terms of $A$ and BB— is not. This nonsmoothness is a major obstacle to applying classic convergence proof techniques of gradient methods to LoRA.
As a result, current theoretical analyses of LoRA fall into one of three categories. First, there are several works that examine LoRA in infinite regimes, such as its asymptotic convergence or its behavior on infinitely wide neural networks [kim2025lora, jang2024lora, zeng2024the, yaras2024compressible, zhang2024riemannianpreconditionedlorafinetuning, hayou2024loraefficientlowrank]. However, these works do not establish a non-asymptotic convergence rate for finite models. Second, many works propose and analyze memory-efficient algorithms that resemble but do not address the original LoRA algorithm. These include GaLore [zhao2024galore] and LoRA variants that only update a single adapter matrix at a time [malinovsky2024randomizedasymmetricchainlora, sokolov2025bernoulliloratheoreticalframeworkrandomized]. Third, some recent works analyze algorithmic frameworks that do extend to LoRA, and do derive a convergence rate [jiang2024a, sun2024improving, ghiasvand2025decentralizedlowrankfinetuninglarge]. However, these works require a strong boundedness assumption that artificially enforces Lipschitz smoothness within the new parametrization. Typically, they require that the adapter matrices $A$ and $B$ are uniformly upper bounded by some constants which appear in the final convergence bound. Considering the existing literature, we seek to answer the question:
How fast does the original LoRA algorithm converge?
In this work, we close the above research gap by establishing the first non-asymptotic convergence analysis of the LoRA gradient descent algorithm, without requiring asynchronous updates, bounded adapter matrices, or Lipschitz smoothness of the reparametrized loss function. We only assume that the original loss function is Lipschitz smooth. Our proof approach involves stacking $B$ and $A$ into a single matrix $V$ and deriving a modified descent lemma for gradient descent with respect to $V$. We show that to achieve descent at each step, the algorithm requires a step size that is “small enough” with respect to the norm of the current parameters and gradient. This recursive relationship between the parameters, gradient, and step size introduces instability in the convergence, resulting in an $O(\frac{1}{\log T})$ convergence rate to a stationary point. If we further assume that the adapter norms are bounded, we can recover the classic $O(\frac{1}{T})$ convergence rate of gradient descent on Lipschitz smooth functions.
Our theoretical analysis reveals two key insights. First, the convergence rate of LoRA does not (and should not) depend on the chosen rank, since gradient descent is a dimension-free algorithm. Second, without additional structural assumptions, the point of convergence of LoRA can be arbitrarily far away from that of full-rank gradient descent. This supports empirical observations that LoRA can yield completely different solutions than full-rank finetuning [shuttleworth2025loravsfinetuningillusion, biderman2024lora].
In Section 2, we review the related work in detail and contextualize our contribution. In Section 3, we provide our convergence analysis, establishing preliminaries in Section 3.1 and providing our main results in Section 3.2. Finally, in Section 4, we discuss the insights from our work.
2 Related Work
The LoRA algorithm, first proposed in [hu2022lora], is motivated by the idea of a viable “low-rank representation” of deep neural networks [oymaklowrank, li2018measuring]. In particular, [aghajanyan-etal-2021-intrinsic] argues that large overparametrized models reside on a lower “intrinsic dimension,” and fine-tuning with a low dimension reparametrization, such as a random subset of the weight matrices, can yield strong performance while minimizing computational expenses.
LoRA is simple, popular, and empirically effective for finetuning large language models. Many variants have been proposed to improve the initialization, performance, memory efficiency, or privacy of the algorithm [shen2025kroneckerlorahybridkroneckerloraadapters, büyükakyüz2024oloraorthonormallowrankadaptation, li2025flatlora, meng2024pissa, wang2024loraga], including but not limited to federated LoRA [10.24963/ijcai.2025/1196], quantized LoRA (QLoRA) [dettmersqlora], and ReLoRA [lialin2024relora]. While LoRA performs well in practice, its behavior in comparison to full-rank training is poorly understood, with empirical studies showing that the two approaches can produce completely different solutions [shuttleworth2025loravsfinetuningillusion, biderman2024lora]. Consequently, the theoretic properties of LoRA are highly relevant for understanding the strengths and weakness of this method.
As highlighted in the introduction, the theory of LoRA convergence has been tackled from many angles. First, many works examine the behavior of LoRA in infinite regimes, whether by characterizing the kinds of solutions LoRA converges to at infinity or analyzing its convergence in the neural tangent kernel (NTK) regime, which defines the linearized training dynamics of an infinitely wide neural network [pmlr-v202-malladi23alazy]. For example, [kim2025lora] argues that LoRA converges to a global minimum with high probability, as long as such a low-rank minimizer exists. The work [zeng2024the] studies the expressive power of the low-rank solutions achieved by LoRA. The work [jang2024lora] shows that LoRA has no spurious local minima in the NTK regime. The LoRA+ [hayou2024loraefficientlowrank] algorithm, which updates the adapter matrices with differently scaled learning rates, is motivated by an NTK convergence analysis. Finally, the gradient flow dynamics of LoRA have been studied specifically for matrix factorization [xu2025understandingGF]. These works do not establish a convergence rate for discrete-time LoRA for general functions.
Second, there are memory-efficient LoRA-like algorithms with concrete convergence analyses that do not apply to the original LoRA formulation. These include GaLore, an algorithm which projects the gradient into a low-rank subspace but maintains a full-rank update of the parameters [zhao2024galore], LDAdam [robert2025ldadam], which performs adaptive optimization steps within lower dimensional subspaces, and Randomized Subspace Optimization (RSO) [chen2025arso]. The LoRA variant LoRA-One incorporates information from the full gradient at initialization [zhang2025loraone], and it is analyzed only for Gaussian data and mean-squared loss. Finally, the variants RAC-LoRA [malinovsky2024randomizedasymmetricchainlora] and Bernoulli-LoRA [sokolov2025bernoulliloratheoreticalframeworkrandomized] only train one adapter matrix while freezing the other as a fixed projection matrix, essentially preserving Lipschitz smoothness while failing to capture the complex training dynamics of actual LoRA implementations. In practice, the LoRA matrices are updated simultaneously, causing nonlinearity and nonsmoothness in the optimized variables. As the original LoRA formulation is widely used in practice, these analyses do not address realistic implementations.
Third, several works prove the convergence of LoRA under a strong boundedness assumption that artificially enforces Lipschitz smoothness. This includes federated LoRA algorithms [sun2024improving, park2025communicationefficientfederatedlowrankupdate, ghiasvand2025decentralizedlowrankfinetuninglarge], for which single-device LoRA is a special case. These works assume that the adapter matrix norms are uniformly bounded by constants. In addition, [jiang2024a] analyzes convergence of LoRA and representation fine-tuning (ReFT) within the same framework, assuming that the singular values of $A$ and $B$ are uniformly upper bounded, which upper bounds the matrix norms by rank-dependent constants that appear in the final convergence result. They also require that the gradient is uniformly bounded. These assumptions force Lipschitz smoothness in terms of the constants and are overly restrictive.
Our work closes the gap in the existing literature by deriving the explicit convergence rate of the original LoRA algorithm for general models, without requiring bounded parameter norms or Lipschitz smoothness of the reparametrized loss function.
3 Convergence Analysis
3.1 Preliminaries
For real-valued matrices $M,N\in\mathbb{R}^{m\times n}$, we denote the Frobenius inner product $\langle\cdot,\cdot\rangle_{F}$ as
$\langle M,N\rangle_{F}=\sum_{i,j}m_{ij}n_{ij}=Tr(M^{T}N),$
and the Frobenius matrix norm $\lVert\cdot\rVert_{F}$ as
$\lVert M\rVert_{F}=\sqrt{\sum_{i,j}|m_{ij}|^{2}}=\sqrt{Tr(M^{T}M)}.$
The Frobenius inner product and norm for matrices are analogous to the Euclidean inner product and norm for vectors. To simplify notation, we also use $\lVert\cdot\rVert$ to denote the Frobenius matrix norm.
We can characterize fine-tuning as the model-agnostic minimization problem
$\min_{W\in\mathbb{R}^{m\times n}}\ell(W_{0}+W),$ (1)
where $\ell:\mathbb{R}^{m\times n}\to\mathbb{R}$ represents the loss function, $W_{0}\in\mathbb{R}^{m\times n}$ represents a frozen (pretrained) weight matrix, and $W\in\mathbb{R}^{m\times n}$ (sometimes denoted as $\Delta W$) represents the update to the frozen weights after fine-tuning. In the LoRA algorithm, we parametrize $W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are the low-rank adapter matrices with rank $r<\min\{m,n\}$. During finetuning, only $A$ and $B$ are optimized while $W_{0}$ remains fixed. To simplify notation, we can reformulate (1) with the function $\mathcal{L}(W)=\ell(W_{0}+W)$ to obtain the LoRA minimization problem,
$\min_{B\in\mathbb{R}^{m\times r},A\in\mathbb{R}^{r\times n}}\mathcal{L}(BA).$ (2)
We herein call $\mathcal{L}:\mathbb{R}^{m\times n}\to\mathbb{R}$ the original loss function. While any kind of optimization can be applied to (2), in this work we consider the prevailing case of LoRA gradient descent, where at time step $t$ the matrices $A$, $B$ are updated simultaneously as follows,
$A_{t+1}=A_{t}-\eta_{t}\nabla_{A}\mathcal{L}(B_{t}A_{t}),\qquad B_{t+1}=B_{t}-\eta_{t}\nabla_{B}\mathcal{L}(B_{t}A_{t}).$ (3)
We require the following standard assumptions for our analysis.
Assumption 1 (Lipschitz smoothness).
The original loss function $\mathcal{L}:\mathbb{R}^{m\times n}\!\to\!\mathbb{R}$ is differentiable, and there exists a constant $L\geq 1$ such that for all $W,W^{\prime}\in\mathbb{R}^{m\times n}$,
$\big\|\nabla\mathcal{L}(W)-\nabla\mathcal{L}(W^{\prime})\big\|_{F}\;\leq\;L\,\big\|W-W^{\prime}\big\|_{F}.$ (4)
Equivalently, $\mathcal{L}$ satisfies the following “descent lemma”
$\mathcal{L}(W^{\prime})\leq\mathcal{L}(W)+\langle\nabla\mathcal{L}(W),W^{\prime}-W\rangle_{F}+\frac{L}{2}\lVert W-W^{\prime}\rVert_{F}^{2}.$ (5)
Assumption 2.
The original loss function $\mathcal{L}$ is lower bounded by a constant $\mathcal{L}^{*}$ such that for all $W\in\mathbb{R}^{m\times n}$, $\mathcal{L}(W)\geq\mathcal{L}^{*}$.
Assumptions 1 and 2 are fundamental to the study of gradient descent convergence. Critically, even with these assumptions, $\mathcal{L}$ is not Lipschitz smooth in $B$ or AA, preventing the analysis of (3) via classic optimization techniques.
3.2 Results
In the following, we detail three steps for proving convergence of LoRA. First, we reformulate the problem (2) as optimization over a single variable $V$ that contains the stacked matrices $B$ and $A^{T}$. We can rewrite the loss function in terms of the outer product $VV^{T}$. Second, we derive a modified “descent lemma,” analogous to the descent lemma for Lipschitz smooth functions (5), that enables descent in one step as long as the step size is small enough with respect to the parameter norm $\lVert V\rVert$ and gradient norm. The last step of the proof is ensuring that the step size does not decrease too quickly, thereby guaranteeing convergence.
Step 1: Restructure problem into $VV^{T}$ form.
We stack the adapter matrices $A^{T}$, $B$, into a single variable $V\in\mathbb{R}^{(m+n)\times r}$. Then the outer product $VV^{T}$ contains $BA$ as follows,
$V=\begin{bmatrix}B\\ A^{T}\end{bmatrix},\qquad VV^{T}=\begin{bmatrix}BB^{T}&BA\\ A^{T}B^{T}&A^{T}A\end{bmatrix}$.
To recover $BA$, we need to extract the top right block from $VV^{T}$. Let $I_{n}$ represent the $n\times n$ identity matrix. Let $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$ represent the “extractor” matrix that extracts the top $m$ rows from a $(m+n)\times(m+n)$ matrix when applied from the left, and let $E_{2}\penalty 10000\ =\penalty 10000\ \begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$ represent the matrix that extracts the right $n$ columns when applied from the right. Then we have that
$BA=E_{1}VV^{T}E_{2}.$
We next define a new function $\mathcal{J}\penalty 10000\ :\penalty 10000\ \mathbb{R}^{(m+n)\times r}\to\mathbb{R}$ such that
$\mathcal{J}(V)=\mathcal{L}(E_{1}VV^{T}E_{2})=\mathcal{L}(BA).$ (6)
By construction, LoRA gradient descent (3) is equivalent to gradient descent on $\mathcal{J}(V)$,
$V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t}),$
since
$\nabla\mathcal{J}(V)=$ $\nabla_{V}[\mathcal{L}(E_{1}VV^{T}E_{2})]$
$=$ $\begin{bmatrix}\nabla_{B}\mathcal{L}(BA)\\ \nabla_{A^{T}}\mathcal{L}(BA)\end{bmatrix}=\begin{bmatrix}\nabla_{B}\mathcal{L}(BA)\\ (\nabla_{A}\mathcal{L}(BA))^{T}\end{bmatrix}$.
Step 2: “Lipschitz-like” descent lemma.
We consider the gradient $\nabla\mathcal{J}(V)\in\mathbb{R}^{(m+n)\times r}$. Let $g(V)=E_{1}VV^{T}E_{2}$, where $g:\mathbb{R}^{(m+n)\times r}\to\mathbb{R}^{m\times n}$. Then $\mathcal{J}(V)=\mathcal{L}(g(V))$. Let $G=\nabla\mathcal{L}(X)|_{X=E_{1}VV^{T}E_{2}}$ represent the gradient of $\mathcal{L}$ evaluated at $E_{1}VV^{T}E_{2}$. Then we have by the chain rule,
$d\mathcal{J}=$ $\langle G,dg\rangle_{F}$
$=$ $\langle G,E_{1}(dVV^{T}+VdV^{T})E_{2}\rangle_{F}$
$=$ $Tr(G^{T}E_{1}(dVV^{T}+VdV^{T})E_{2})$
$=$ $Tr(G^{T}E_{1}dVV^{T}E_{2})+Tr(G^{T}E_{1}VdV^{T}E_{2})$
$=$ $Tr(V^{T}E_{2}G^{T}E_{1}dV)+Tr(dV^{T}E_{2}G^{T}E_{1}V)$
$=$ $\langle E_{1}^{T}GE_{2}^{T}V,dV\rangle_{F}+\langle E_{2}G^{T}E_{1}V,dV\rangle_{F}$
$=$ $\langle 2Sym(E_{1}^{T}GE_{2}^{T})V,dV\rangle_{F},$
where $Sym(A)=\frac{A+A^{T}}{2}$. So we have
$\nabla\mathcal{J}(V)=2Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}VV^{T}E_{2})E_{2}^{T})V.$ (7)
The function $\mathcal{J}$ is not Lipschitz smooth, due to the multiplicative factor of $V$ in $\nabla\mathcal{J}(V)$. However, we can still obtain a “descent lemma” (Lemma 1) that is analogous to the classic descent lemma of Lipschitz smooth functions (5), but with more higher-order terms.
Lemma 1.
For $\mathcal{J}(V)$ defined in (6), we have for all $V_{1},V_{2}\in\mathbb{R}^{(m+n)\times r}$,
$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{\sqrt{2}L}{3}\lVert V_{2}-V_{1}\rVert^{3}$
$+\frac{\sqrt{2}L}{4}\lVert V_{2}-V_{1}\rVert^{4}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
Proof. See Appendix A.2.
Step 3: Control step size to obtain convergence.
Based on Lemma 1, we can pick $\eta_{t}$ small enough to minimize the higher-order terms, thereby guaranteeing descent in one step as stated next.
Lemma 2.
For any $V_{t}\in\mathbb{R}^{(m+n)\times r}$ and $V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, suppose that
$\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\}.$ (8)
Then we have
$\mathcal{J}(V_{t+1})\leq\mathcal{J}(V_{t})-\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}.$ (9)
Proof. See Appendix A.3.
Lemma 2 states that to update $V_{t+1}$, we require the norm of the previous iterate $\lVert V_{t}\rVert$ and the gradient of the original loss function $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$. By summing (9) over all t=0,,T1t=0,...,T-1 and telescoping sums, we determine that the minimum gradient norm over $t$ is upper bounded as
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\sum_{t=0}^{T-1}\eta_{t}}.$ (10)
However, (10) is not sufficient to conclude convergence; we also need to show that the sum $\sum_{t=0}^{T-1}\eta_{t}$ diverges as $T$ goes to infinity rather than converging to a finite value. Based on (8), this automatically holds if $\lVert V_{t}\rVert^{2}$ is uniformly upper bounded, leading to $\sum_{t=0}^{T-1}\eta_{t}=\Theta(T)$. However, if the parameter norm is not bounded, then $\lVert V_{t}\rVert^{2}$ may grow to infinity. In the proof of Theorem 3, we show that even in the worst case, we have $\lVert V_{t}\rVert^{2}=O(t)$. The sum $\sum_{t=0}^{T-1}\eta_{t}$ is therefore lower bounded by a harmonic series which is $\Theta(\log(T))$, leading to Theorem 3.
Theorem 3.
Suppose Assumption 1 and 2 hold. We perform $T$ steps of LoRA gradient descent (3), where at time step $t$ the step size $\eta_{t}$ is determined as
$\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\}.$
Then after $T$ steps, we have
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}=O\left(\frac{1}{\log T}\right),$ (11)
where the $O(\cdot)$ notation hides dependence on $V_{0}$, $\mathcal{J}(V_{0})$, $\mathcal{L}^{*}$, and $L$. Moreover, if there exists a constant $C>0$ such that for all $V_{t}$, $\lVert V_{t}\rVert\leq C$, then
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}=O\left(\frac{1}{T}\right),$ (12)
where the $O(\cdot)$ notation hides dependence on $\mathcal{J}(V_{0})$, $\mathcal{L}^{*}$, $L$, and $C$.
Proof. See Appendix A.4.
Theorem 3 states that LoRA gradient descent converges to a stationary point at $O(\frac{1}{\log T})$ rate. If $\lVert V_{t}\rVert$ is uniformly bounded, the rate improves to $O(1/T)$. This matches the standard convergence rate of gradient descent on Lipschitz smooth functions and the rates derived by existing works that analyze LoRA under the bounded parameter assumption, which essentially forces $\mathcal{J}$ to be Lipschitz smooth. Our results demonstrate that the behavior of LoRA is governed by complex interactions between $\eta_{t}$, $\lVert V_{t}\rVert$, and $\lVert\nabla\mathcal{J}(V_{t})\rVert$, which slows convergence. However, even if $\lVert V_{t}\rVert$ grows and forces $\eta_{t}$ to decrease at each iteration, the algorithm still converges; a simple explanation is that if $\lVert V_{t}\rVert$ grows, $\eta_{t}$ decreases, which slows the growth of $\lVert V_{t}\rVert$.
4 Discussion
In this work, we show for first time that the original LoRA gradient descent algorithm achieves a convergence rate of $O(\frac{1}{\log T})$, without requiring bounded adapter matrices or Lipschitz smoothness of the reparametrized loss function.
Our results yield several interesting insights into the behavior of LoRA. First, the choice of extractor matrices $E_{1}$, $E_{2}$, has little to no impact on the proof, suggesting that other subsets of the $VV^{T}$ matrix could be extracted as alternate parameter-efficient reparametrizations to yield similar results. Moreover, the $VV^{T}$ form is exactly the symmetric Burer-Monteiro parametrization [burer_nonlinear_2003], and the results of our analysis can be trivially extended to show that gradient descent with Burer-Monteiro also converges to stationary points for general Lipschitz smooth functions. Second, the convergence rate does not explicitly depend on the chosen LoRA rank $r$, except that it is $O(\frac{1}{\lVert V_{0}\rVert^{2}})$. This is expected because $r$ just controls the dimension of the gradient descent algorithm, which is a dimension-free algorithm. Finally, our results shed light onto the complex relationship between the parameter norm and the LoRA training dynamics, which has been alluded to in past work. In particular, we observe that the convergence of LoRA is slowed if the iterates are progressing away from the origin, and it is accelerated if they are moving towards the origin. This “position dependency” is not typical of gradient algorithms, and may be due to the fact that the LoRA reparametrization creates a stationary point at $V=0$ regardless of the structure of the original function. Consequently, LoRA may converge to the origin even if the original global minima is arbitrarily far away, explaining the divergence in behavior between LoRA and full-rank fine-tuning.
Future research directions may include determining convergence rates on convex or strongly convex loss functions, whose geometric properties may change under the LoRA reparametrization, as well as analyzing the stochastic gradient descent setting. A major challenge of this setting is deriving a fourth-moment bound on the noise.
Appendix A Proofs
A.1 Helper Lemmas
Lemma 4.
(Weighted) AM-GM inequality. Given the nonnegative values $x_{1},x_{2},w_{1},w_{2}$, we have
$\frac{w_{1}x_{1}+w_{2}x_{2}}{w_{1}+w_{2}}\geq(x_{1}^{w_{1}}x_{2}^{w_{2}})^{\frac{1}{w_{1}+w_{2}}}.$ (13)
In particular, for $w_{1}=w_{2}=1$, we have
$x_{1}+x_{2}\geq 2\sqrt{x_{1}x_{2}}.$ (14)
For the proof of Lemma 4, see Section 2.5 of [hardy_inequalities_1934].
Lemma 5.
For any matrix $A\in\mathbb{R}^{(m+n)\times(m+n)}$ and the extractor matrices $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$, and $E_{2}=\begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$
$\lVert E_{1}AE_{2}\rVert\leq\lVert A\rVert.$
Moreover, for any matrix $B\in\mathbb{R}^{m\times n}$,
$\lVert E_{1}^{T}BE_{2}^{T}\rVert=\lVert B\rVert.$
Finally, if $A\in\mathbb{R}^{(m+n)\times(m+n)}$ is symmetric,
$\lVert E_{1}AE_{2}\rVert\leq\frac{1}{\sqrt{2}}\lVert A\rVert.$ (15)
Proof.
The first statement is true because $E_{1}AE_{2}$ contains a subset of the entries of $A$. The second statement is true because $E_{1}^{T}BE_{2}^{T}$ produces a matrix that only contains the entries of B and zeros. The last statement can be shown by observing that $E_{1}AE_{2}$ extracts the upper right $m\times n$ block from $A$, specifically the elements at $i,j$ where $1\leq i\leq m$ and $m+1\leq j\leq m+n$. So if we have
$A=\begin{bmatrix}A_{11}&A_{12}\\ A_{12}^{T}&A_{22}\end{bmatrix}$,
Then $\lVert A\rVert^{2}=\lVert A_{11}\rVert^{2}+\lVert A_{22}\rVert^{2}+2\lVert A_{12}\rVert^{2}$ and $\lVert A_{12}\rVert\leq\frac{1}{\sqrt{2}}\lVert A\rVert$. ∎
Lemma 6.
Let $\{a_{n}\}_{n\geq 0}$ be a real, nonnegative sequence, and let $c>0$ be a fixed constant. If the series $\sum_{n=1}^{\infty}\min(a_{n},c)$ converges, then the series $\sum_{n=1}^{\infty}a_{n}$ also converges.
Proof.
Let $b_{n}=\min(a_{n},c)$. Since $\sum_{n=1}^{\infty}b_{n}$ converges, the terms $b_{n}$ must go to zero, so there exists $N>0$ such that for $n\geq N$,
$b_{n}\leq\frac{c}{2}.$
This also implies that for $n\geq N$, $a_{n}\leq\frac{c}{2}<c$. So we have that $\sum_{n=N}^{\infty}a_{n}=\sum_{n=N}^{\infty}b_{n}$, and both these tail sums must converge. Then our original series can be written in terms of a finite sum and the tail sum, both of which are finite.
$\sum_{n=1}^{\infty}a_{n}=\sum_{n=1}^{N-1}a_{n}+\sum_{n=N}^{\infty}a_{n}=\sum_{n=1}^{N-1}a_{n}+\sum_{n=N}^{\infty}b_{n}.$
A.2 Proof of Lemma 1
Proof.
Based on the expression for the gradient $\nabla\mathcal{J}(V)$ in (7), we have
$\nabla\mathcal{J}(V_{2})-\nabla\mathcal{J}(V_{1})=$ $2\big[Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{2}V_{2}^{T}E_{2})E_{2}^{T})V_{2}-Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})V_{1}\big],$
$=$ $2\big[Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{2}V_{2}^{T}E_{2})E_{2}^{T})V_{2}-Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})V_{2}$
$+Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})V_{2}-Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})V_{1}\big],$
$=$ $2\big[(Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{2}V_{2}^{T}E_{2})E_{2}^{T})-Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T}))V_{2}$
$+Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})(V_{2}-V_{1})\big],$
$=$ $2\big[Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{2}V_{2}^{T}E_{2})E_{2}^{T}-E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})V_{2}$
$+Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})(V_{2}-V_{1})\big],$
$=$ $2\big[Sym(E_{1}^{T}(\nabla\mathcal{L}(E_{1}V_{2}V_{2}^{T}E_{2})-\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2}))E_{2}^{T})V_{2}$
$+Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})(V_{2}-V_{1})\big].$ (16)
If we denote $d=V_{2}-V_{1}$, and parametrize $\mathcal{J}$ between $V_{1}$ and $V_{2}$ such that for $t\in[0,1]$, then
$\phi(t)=$ $\mathcal{J}(V_{1}+td),$
$\phi^{\prime}(t)=$ $\langle\nabla\mathcal{J}(V_{1}+td),d\rangle_{F}.$
By the fundamental theorem of calculus, we have
$\mathcal{J}(V_{2})-\mathcal{J}(V_{1})$ $=\phi(1)-\phi(0)=\int_{0}^{1}\phi^{\prime}(t)dt,$
$=\int_{0}^{1}\langle\nabla\mathcal{J}(V_{1}),d\rangle_{F}\;dt+\int_{0}^{1}\langle\nabla\mathcal{J}(V_{1}+td)-\nabla\mathcal{J}(V_{1}),d\rangle_{F}\;dt,$
$=\langle\nabla\mathcal{J}(V_{1}),d\rangle_{F}+\int_{0}^{1}\langle\nabla\mathcal{J}(V_{1}+td)-\nabla\mathcal{J}(V_{1}),d\rangle_{F}\;dt,$
$\leq\langle\nabla\mathcal{J}(V_{1}),d\rangle+\int_{0}^{1}\lVert\nabla\mathcal{J}(V_{1}+td)-\nabla\mathcal{J}(V_{1})\rVert\lVert d\rVert dt,$ (17)
where in the last step we use the Cauchy-Schwarz inequality and the fact that integrals preserve inequalities. From (16), we have
$\lVert\nabla\mathcal{J}(V_{1}+td)-\nabla\mathcal{J}(V_{1})\rVert=$ $2\lVert Sym(E_{1}^{T}(\nabla\mathcal{L}(E_{1}(V_{1}+td)(V_{1}+td)^{T}E_{2})-\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2}))E_{2}^{T})(V_{1}+td)$
$+Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})td\rVert,$
$\overset{\text{Cauchy-Schwarz}}{\leq}$ $2\lVert Sym(E_{1}^{T}(\nabla\mathcal{L}(E_{1}(V_{1}+td)(V_{1}+td)^{T}E_{2})-\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2}))E_{2}^{T})\rVert\lVert V_{1}+td\rVert$
$+2\lVert Sym(E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T})\rVert\lVert td\rVert,$
$\overset{\lVert Sym(A)\rVert\leq\lVert A\rVert}{\leq}$ $2\lVert E_{1}^{T}(\nabla\mathcal{L}(E_{1}(V_{1}+td)(V_{1}+td)^{T}E_{2})-\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2}))E_{2}^{T}\lVert\rVert V_{1}+td\rVert$
$+2\lVert E_{1}^{T}\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})E_{2}^{T}\rVert\lVert td\rVert,$
[…]
$+2t\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert,$
[…]
$=$ $\sqrt{2}L\lVert(V_{1}+td)(td)^{T}+tdV_{1}^{T}\lVert\rVert V_{1}+td\rVert+2t\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert,$
$\leq$ $\sqrt{2}Lt\lVert d\rVert\lVert V_{1}+td\rVert^{2}+\sqrt{2}Lt\lVert d\rVert\lVert V_{1}\rVert\lVert V_{1}+td\rVert+2t\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert,$
$\leq$ $\sqrt{2}Lt\lVert d\rVert\lVert V_{1}+td\rVert^{2}+\sqrt{2}Lt\lVert d\rVert\lVert V_{1}\rVert^{2}+\sqrt{2}Lt^{2}\lVert d\rVert^{2}+2t\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert.$
We plug this expression into the integral in (17) and pull out terms that do not depend on $t$ to obtain
$\int_{0}^{1}\lVert\nabla\mathcal{J}(V_{1}+td)-\nabla\mathcal{J}(V_{1})\rVert\lVert d\rVert dt\leq$ $\sqrt{2}L\lVert d\rVert^{2}\int_{0}^{1}t\lVert V_{1}+td\rVert^{2}dt+\sqrt{2}L\lVert d\rVert^{2}\lVert V_{1}\rVert^{2}\int_{0}^{1}tdt$
$+\sqrt{2}L\lVert d\rVert^{3}\int_{0}^{1}t^{2}dt+2\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert^{2}\int_{0}^{1}tdt,$
$=$ $\sqrt{2}L\lVert d\rVert^{2}(\frac{1}{2}\lVert V_{1}\rVert^{2}+\frac{2}{3}\langle V_{1},d\rangle_{F}+\frac{1}{4}\lVert d\rVert^{2})+\frac{\sqrt{2}}{2}L\lVert d\rVert^{2}\lVert V_{1}\rVert^{2}$
$+L\frac{\sqrt{2}}{3}\lVert d\rVert^{3}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert^{2},$
$\overset{\text{Cauchy-Schwarz}}{\leq}$ $\frac{\sqrt{2}}{2}L\lVert d\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{2\sqrt{2}}{3}L\lVert d\rVert^{3}\lVert V_{1}\rVert+\frac{\sqrt{2}}{4}L\lVert d\rVert^{4}+\frac{\sqrt{2}}{2}L\lVert d\rVert^{2}\lVert V_{1}\rVert^{2}$
$+L\frac{\sqrt{2}}{3}\lVert d\rVert^{3}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert^{2},$
$=$ $\frac{2\sqrt{2}}{3}L\lVert d\rVert^{3}\lVert V_{1}\rVert+\frac{\sqrt{2}}{4}L\lVert d\rVert^{4}+\sqrt{2}L\lVert d\rVert^{2}\lVert V_{1}\rVert^{2}+L\frac{\sqrt{2}}{3}\lVert d\rVert^{3}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert d\rVert^{2}.$
We therefore have for all $V_{1}$, $V_{2}$,
$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+L\frac{\sqrt{2}}{3}\lVert V_{2}-V_{1}\rVert^{3}$
$+\frac{\sqrt{2}}{4}L\lVert V_{2}-V_{1}\rVert^{4}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
A.3 Proof of Lemma 2
Proof.
We apply the descent lemma (Lemma 1) with $V_{1}=V_{t}$, $V_{2}=V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, which yields
$\mathcal{J}(V_{t+1})\leq$ $\mathcal{J}(V_{t})-\eta_{t}\langle\nabla\mathcal{J}(V_{t}),\nabla\mathcal{J}(V_{t})\rangle_{F}+\frac{2\sqrt{2}}{3}L\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}\lVert V_{t}\rVert+\sqrt{2}L\eta_{t}^{2}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\lVert V_{t}\rVert^{2}+\frac{\sqrt{2}L}{3}\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}$
$+\frac{\sqrt{2}}{4}\eta_{t}^{4}L\lVert\nabla\mathcal{J}(V_{t})\rVert^{4}+\eta_{t}^{2}\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$=$ $\mathcal{J}(V_{t})-\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}+\frac{2\sqrt{2}}{3}L\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}\lVert V_{t}\rVert+\eta_{t}^{2}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}(\sqrt{2}L\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)$
$+\frac{\sqrt{2}L}{3}\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}+\frac{\sqrt{2}L}{4}\eta_{t}^{4}\lVert\nabla\mathcal{J}(V_{t})\rVert^{4}.$
We want to select $\eta_{t}$ to minimize the last four terms, such that they sum to a value smaller than $\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}$. This will guarantee descent in function value $\mathcal{J}$. Let
$\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\}.$
In the following, we prove that $\eta_{t}$ is smaller than various upper bounds that will allow us to achieve a clean descent result. First, since $L\geq 1$, we have the following bound,
$\eta_{t}\leq\frac{1}{5(\sqrt{2}L\lVert V_{t}\rVert^{2}+\sqrt{2}L\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)}\leq\frac{1}{5(\sqrt{2}L\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)}.$ (18)
For the next bound, we have from Cauchy-Schwarz and (7) that
$\lVert\nabla\mathcal{J}(V)\rVert\leq 2\lVert\nabla\mathcal{L}(E_{1}VV^{T}E_{2})\rVert\lVert V\rVert.$ (19)
Therefore, from (14) of Lemma 4, we have $\frac{1}{a+b}\leq\frac{1}{2\sqrt{ab}}$ for $a,b>0$, such that $\eta_{t}$ satisfies the following bound,
$\eta_{t}\leq\frac{1}{5\sqrt{2}L\lVert V_{t}\rVert^{2}+5\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}\leq$ $\left(\frac{1}{100\sqrt{2}L\lVert V_{t}\rVert^{2}\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}\right)^{1/2},$
[…]
$\leq$ $\left(\frac{3}{10\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert\lVert V_{t}\rVert}\right)^{1/2}.$ (21)
For the next bound, by (13) of Lemma 4 with $w_{1}=1$, $w_{2}=2$, $x_{1}=5\sqrt{2}\lVert V_{t}\rVert^{2}$, $x_{2}=\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$, we have
$5\sqrt{2}L\lVert V_{t}\rVert^{2}+2\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert\geq$ $3\left(5\sqrt{2}L\lVert V_{t}\rVert^{2}\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert^{2}\right)^{1/3},$
[…]
So $\eta_{t}$ also satisfies the bound
$\eta_{t}\leq$ $\frac{1}{5\sqrt{2}L\lVert V_{t}\rVert^{2}+5\sqrt{2}L\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert},$
$\leq$ $\frac{1}{5\sqrt{2}L\lVert V_{t}\rVert^{2}+2\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert},$
$\leq$ $\left(\frac{4}{135\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}}\right)^{1/3},$
$\leq$ $\left(\frac{4}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}}\right)^{1/3}.$ (22)
Finally, we also want to show
$\eta_{t}\leq\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2}.$ (23)
Let us assume that $\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2}<1$, because if this is not true, then the result is trivial. We can break down the casework as follows.
  • Case 1: $\lVert V_{t}\rVert=0$. In this case, we have by (7) that $\lVert\nabla\mathcal{J}(V_{t})\rVert=0$. So the result is trivially shown.
  • Case 2: $0<\lVert V_{t}\rVert\leq 1$. In this case, we have from (19)
    $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert\geq\frac{\lVert\nabla\mathcal{J}(V_{t})\rVert}{2\lVert V_{t}\rVert}.$
    Then we have
    $\eta_{t}\leq$ $\frac{1}{5\sqrt{2}L\lVert V_{t}\rVert^{2}+5L\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert},$
    $\leq$ $\frac{1}{5\sqrt{2}L\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert},$
    $\leq$ $\frac{2\lVert V_{t}\rVert}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert},$
    $\leq$ $\frac{2}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert},$
    $\leq$ $\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\leq\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2},$
    where in the last step we use the fact that $\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2}<1$.
  • Case 3: $\lVert V_{t}\rVert>1$. From (20), we have
    $\eta_{t}\leq$ $\left(\frac{1}{50\sqrt{2}L\lVert V_{t}\rVert\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2},$
    $\leq$ $\left(\frac{1}{50\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2},$
    $\leq$ $\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2}.$
In conclusion, with the choice of $\eta_{t}$ in (8), we can substitute the derived bounds on $\eta_{t}$ into Lemma 1, achieving descent in one step.
[…]
$\mathcal{J}(V_{t+1})\leq$ $\mathcal{J}(V_{t})-\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}+\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}+\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}+\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}+\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}$
$\leq$ $\mathcal{J}(V_{t})-\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}.$
A.4 Proof of Theorem 3
Proof.
From Lemma 2, we know that if
$\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\},$
then we have descent in one step
$\mathcal{J}(V_{t+1})-\mathcal{J}(V_{t})\leq$ $-\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}.$
Rearranging this equation and summing on both sides yields
$\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\mathcal{J}(V_{t})-\mathcal{J}(V_{t+1}),$
$\sum_{t=0}^{T-1}\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\sum_{t=0}^{T-1}\mathcal{J}(V_{t})-\mathcal{J}(V_{t+1}),$
$=$ $\mathcal{J}(V_{0})-\mathcal{J}(V_{T}),$
[…]
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\sum_{t=0}^{T-1}\eta_{t}\leq$ $5(\mathcal{J}(V_{0})-\mathcal{L}^{*}),$
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\sum_{t=0}^{T-1}\eta_{t}}.$
So to show convergence to a stationary point, we need to show that the sum of learning rates $\sum_{t=0}^{\infty}\eta_{t}$ diverges. From Lipschitz smoothness of $\mathcal{L}$, we have for all $W\in\mathbb{R}^{m\times n}$ (See Lemma 2.28 in [garrigos2024handbookconvergencetheoremsstochastic]),
$\lVert\nabla\mathcal{L}(W)\rVert^{2}\leq$ $2L(\mathcal{L}(W)-\mathcal{L}^{*})$
$\lVert\nabla\mathcal{L}(W)\rVert\leq$ $(2L(\mathcal{L}(W)-\mathcal{L}^{*}))^{1/2}$
$\lVert\nabla\mathcal{L}(E_{1}VV^{T}E_{2})\rVert\leq$ $(2L(\mathcal{L}(E_{1}VV^{T}E_{2})-\mathcal{L}^{*}))^{1/2}=(2L(\mathcal{J}(V)-\mathcal{L}^{*}))^{1/2}.$
We can obtain the following lower bound, using the fact that $\mathcal{J}(V_{t})$ is decreasing with $t$,
$\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)}\geq$ $\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+(2L(\mathcal{J}(V_{t})-\mathcal{L}^{*}))^{1/2})},$
$\geq$ $\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})}.$
Let $D_{t}=5\sqrt{2}L(\lVert V_{t}\rVert^{2}+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})$, such that we can lower bound $\eta_{t}$ as follows,
$\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\}\geq\min\{\frac{1}{D_{t}},1\}.$
To show the divergence of $\sum_{t=0}^{\infty}\eta_{t}$, it is sufficient to show the divergence of the series $\sum_{t=0}^{\infty}\min\{\frac{1}{D_{t}},1\}$. Moreover, by the contrapositive of Lemma 6, we only need to show divergence of $\sum_{t=0}^{\infty}\frac{1}{D_{t}}$. We first show that we achieve divergence if the adapter norms are bounded. If $\lVert V_{t}\rVert\leq C$ for all $t$, then $\eta_{t}$ is lower bounded by a nonzero constant $\eta>0$
$\eta:=\min\{\frac{1}{5\sqrt{2}L(C^{2}+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})},1\}\leq\eta_{t},$
and we prove (12) of Theorem 3 as follows
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\eta T}=O(1/T).$
Now we consider the general case where the adapter norms are not necessarily bounded. In the following, we show that $D_{t}=O(t)$. We first control the growth of $\lVert V_{t}\rVert^{2}$, showing that it increases at most linearly with $t$. We have
$\lVert V_{t+1}\rVert^{2}=$ $\lVert V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\leq$ $\lVert V_{t}\rVert^{2}+2\eta_{t}|\langle V_{t},\nabla\mathcal{J}(V_{t})\rangle_{F}|+\eta_{t}^{2}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\overset{\eta_{t}\leq 1}{\leq}$ $\lVert V_{t}\rVert^{2}+2\eta_{t}|\langle V_{t},\nabla\mathcal{J}(V_{t})\rangle_{F}|+\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\leq$ $\lVert V_{t}\rVert^{2}+\eta_{t}(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{J}(V_{t})\rVert^{2})+\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\leq$ $\lVert V_{t}\rVert^{2}+\frac{\lVert V_{t}\rVert^{2}}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)}+2\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\leq$ $\lVert V_{t}\rVert^{2}+\frac{1}{5\sqrt{2}L}+2\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\lVert V_{t+1}\rVert^{2}-\lVert V_{t}\rVert^{2}\leq$ $\frac{1}{5\sqrt{2}L}+2\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\sum_{t=0}^{T-1}\lVert V_{t+1}\rVert^{2}-\lVert V_{t}\rVert^{2}\leq$ $\frac{T}{5\sqrt{2}L}+2\sum_{t=0}^{T-1}\eta_{t}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2},$
$\leq$ $\frac{T}{5\sqrt{2}L}+10(\mathcal{J}(V_{0})-\mathcal{L}^{*}),$
$\lVert V_{T}\rVert^{2}\leq$ $\lVert V_{0}\rVert^{2}+\frac{T}{5\sqrt{2}L}+10(\mathcal{J}(V_{0})-\mathcal{L}^{*}).$
We have that $D_{t}$ is upper bounded by a linear term,
$D_{t}=5\sqrt{2}L(\lVert V_{t}\rVert^{2}+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})\leq t+5\sqrt{2}L(\lVert V_{0}\rVert^{2}+10(\mathcal{J}(V_{0})-\mathcal{L}^{*})+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2}).$
We can therefore lower bound the series with a harmonic series as follows,
$\sum_{t=0}^{T-1}\frac{1}{D_{t}}\geq$ $\sum_{t=0}^{T-1}\frac{1}{t+5\sqrt{2}L(\lVert V_{0}\rVert^{2}+10(\mathcal{J}(V_{0})-\mathcal{L}^{*})+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})},$
$=$ $\Theta(\log(T)).$
Finally,
$\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\sum_{t=0}^{T-1}\eta_{t}}=O\left(\frac{1}{\log T}\right).$

### v2

@@
 In this work, we provide for the first time a non-asymptotic convergence analysis of the original LoRA gradient descent algorithm, which reflects widespread practice, without such assumptions.
-Our work relies on three key steps: i) reformulating the problem in terms of the outer product of the stacked adapter matrices, ii) a modified descent lemma for the “Lipschitz-like” reparametrized function, and iii) controlling the step size.
+Our work relies on three key steps: i) reformulating the problem in terms of the outer product of the stacked adapter matrices, ii) a modified descent lemma for the “Lipschitz-like” reparametrized function, and iii) controlling the learning rate.
 With this approach, we prove that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, where $T$ is the number of iterations.
+We conduct numerical experiments to validate our theoretical findings.
+Machine Learning, ICML
 1 Introduction
 Modern applications of large language models (LLMs) typically involve self-supervised pretraining of a large foundation model, which is later fine-tuned on a smaller task-specific dataset through supervised learning.
-Parameter efficient fine-tuning methods, which only train a small number of additional parameters, can efficiently adapt LLMs to a multitude of downstream tasks while maintaining comparable performance to fine-tuning of the full model parameters.
-One highly popular method is LoRA, or Low-Rank Adaptation [hu2022lora], which trains a low-rank representation of the model parameters.
+Parameter efficient fine-tuning methods, which only train a small number of parameters, can efficiently adapt LLMs to a multitude of downstream tasks while maintaining comparable performance to fine-tuning of the full model parameters.
+One highly popular method is LoRA, or Low-Rank Adaptation (Hu et al., 2022), which trains a low-rank representation of the model parameters.
 Specifically, once the pretraining stage yields a model weight matrix $W_{0}$, LoRA only updates low-rank adapter matrices $A$, $B$, such that the final weights are $W_{0}+BA$.
 The matrices $A$ and $B$ can be updated with any optimization method, including gradient descent.
-However, even this simple algorithm is challenging to analyze; as observed in [sun2024improving] and [malinovsky2024randomizedasymmetricchainlora], even if the original loss function is Lipschitz smooth, the new loss function— reparametrized in terms of $A$ and BB— is not.
-This nonsmoothness is a major obstacle to applying classic convergence proof techniques of gradient methods to LoRA.
+However, even this simple algorithm is challenging to analyze; as observed in (Sun et al., 2024; Malinovsky et al., 2024), even if the original loss function is Lipschitz smooth, the new loss function— reparametrized in terms of $A$ and BB— is not.
+This nonsmoothness is a major obstacle to applying classic convergence analysis techniques to LoRA.
 As a result, current theoretical analyses of LoRA fall into one of three categories.
-First, there are several works that examine LoRA in infinite regimes, such as its asymptotic convergence or its behavior on infinitely wide neural networks [kim2025lora, jang2024lora, zeng2024the, yaras2024compressible, zhang2024riemannianpreconditionedlorafinetuning, hayou2024loraefficientlowrank].
+First, there are several works that examine LoRA in infinite regimes, such as its asymptotic convergence or its behavior on infinitely wide neural networks (Kim et al., 2025; Jang et al., 2024; Zeng and Lee, 2024; Yaras et al., 2024; Zhang and Pilanci, 2024; Hayou et al., 2024).
 However, these works do not establish a non-asymptotic convergence rate for finite models.
 Second, many works propose and analyze memory-efficient algorithms that resemble but do not address the original LoRA algorithm.
-These include GaLore [zhao2024galore] and LoRA variants that only update a single adapter matrix at a time [malinovsky2024randomizedasymmetricchainlora, sokolov2025bernoulliloratheoreticalframeworkrandomized].
-Third, some recent works analyze algorithmic frameworks that do extend to LoRA, and do derive a convergence rate [jiang2024a, sun2024improving, ghiasvand2025decentralizedlowrankfinetuninglarge].
-However, these works require a strong boundedness assumption that artificially enforces Lipschitz smoothness within the new parametrization.
-Typically, they require that the adapter matrices $A$ and $B$ are uniformly upper bounded by some constants which appear in the final convergence bound.
+These include GaLore (Zhao et al., 2024) and LoRA variants that only update a single adapter matrix at a time (Malinovsky et al., 2024; Sokolov et al., 2025).
+Third, some recent works analyze algorithmic frameworks that do extend to LoRA, and do derive a convergence rate (Jiang et al., 2024; Sun et al., 2024; Ghiasvand et al., 2025).
+However, these works require that the adapter matrices $A$ and $B$ are uniformly upper bounded by some constants which appear in the final convergence bound.
+These restrictive assumptions artificially enforce Lipschitz smoothness within the new parametrization, essentially reducing the problem to standard gradient descent with no substantially new proof techniques.
 Considering the existing literature, we seek to answer the question:
 How fast does the original LoRA algorithm converge?
-In this work, we close the above research gap by establishing the first non-asymptotic convergence analysis of the LoRA gradient descent algorithm, without requiring asynchronous updates, bounded adapter matrices, or Lipschitz smoothness of the reparametrized loss function.
-We only assume that the original loss function is Lipschitz smooth.
-Our proof approach involves stacking $B$ and $A$ into a single matrix $V$ and deriving a modified descent lemma for gradient descent with respect to $V$.
-We show that to achieve descent at each step, the algorithm requires a step size that is “small enough” with respect to the norm of the current parameters and gradient.
-This recursive relationship between the parameters, gradient, and step size introduces instability in the convergence, resulting in an $O(\frac{1}{\log T})$ convergence rate to a stationary point.
+In this work, we close the above research gap by establishing the first non-asymptotic convergence analysis of LoRA gradient descent, without requiring asynchronous updates, bounded adapter matrices, or Lipschitz smoothness of the reparametrized loss function.
+We only assume that the original loss function is Lipschitz smooth and lower bounded.
+Our novel proof approach involves stacking $B$ and $A$ into a single matrix $V$ and deriving a modified “Lipschitz-like” descent lemma for gradient descent with respect to $V$.
+We prove that to achieve descent at each step, the algorithm requires a learning rate that is “small enough” with respect to the norm of the current parameters and gradient.
+This recursive relationship between the parameters, gradient, and learning rate introduces instability in the convergence, resulting in an $O(\frac{1}{\log T})$ convergence rate to a stationary point.
 If we further assume that the adapter norms are bounded, we can recover the classic $O(\frac{1}{T})$ convergence rate of gradient descent on Lipschitz smooth functions.
-Our theoretical analysis reveals two key insights.
+By analyzing LoRA under minimal assumptions, we can make rigorous observations about its behavior.
 First, the convergence rate of LoRA does not (and should not) depend on the chosen rank, since gradient descent is a dimension-free algorithm.
-Second, without additional structural assumptions, the point of convergence of LoRA can be arbitrarily far away from that of full-rank gradient descent.
-This supports empirical observations that LoRA can yield completely different solutions than full-rank finetuning [shuttleworth2025loravsfinetuningillusion, biderman2024lora].
+Second, we formalize the complex relationship between the parameter norm and the LoRA training dynamics, which has been alluded to in past work.
+We find that LoRA displays a “position dependency,” in that the convergence is slowed if the iterates are moving away from the origin and accelerated if they are moving towards the origin.
+This is due to the geometric changes introduced by the LoRA reparametrization, including the creation of a stationary point at the origin regardless of the structure of the original loss function.
+Motivated by our theoretical insights, we perform experiments to examine the impact of learning rate choice under the LoRA reparametrization in practice.
+We train logistic regression and ResNet-18 models on the CIFAR-10 images dataset with adaptive and normalized learning rates that directly stem from theory, and we compare their performance against constant learning rates of similar values.
+We find that adjusting the learning rate based on the parameter norm or gradient norm can both accelerate and stabilize training by accounting for the unique structure induced by the low-rank reparametrization.
+Our contributions can be summarized as follows:
+  • We prove for the first time that LoRA gradient descent converges to a stationary point at rate $O(\frac{1}{\log T})$, only assuming that the original loss function is Lipschitz smooth and lower bounded.
+  • Based on our theory, we propose and empirically validate practical methods for LoRA learning rate selection that demonstrate superior convergence.
 In Section 2, we review the related work in detail and contextualize our contribution.
-In Section 3, we provide our convergence analysis, establishing preliminaries in Section 3.1 and providing our main results in Section 3.2.
-Finally, in Section 4, we discuss the insights from our work.
+In Section 3, we provide our convergence analysis, establishing preliminaries in Section 3.1, providing our main results in Section 3.2 and discussion in Section 3.3.
+Finally, in Section 4, we provide the results of our experiments.
+Code is open-sourced at the anonymous Github repository https://anonymous.4open.science/r/lora-C709/.
 2 Related Work
-The LoRA algorithm, first proposed in [hu2022lora], is motivated by the idea of a viable “low-rank representation” of deep neural networks [oymaklowrank, li2018measuring].
-In particular, [aghajanyan-etal-2021-intrinsic] argues that large overparametrized models reside on a lower “intrinsic dimension,” and fine-tuning with a low dimension reparametrization, such as a random subset of the weight matrices, can yield strong performance while minimizing computational expenses.
+The LoRA algorithm, first proposed in (Hu et al., 2022), is motivated by the idea of a viable “low-rank representation” of deep neural networks (Oymak et al., 2019; Li et al., 2018).
+In particular, (Aghajanyan et al., 2021) argues that large overparametrized models reside on a lower “intrinsic dimension,” and fine-tuning with a low dimension reparametrization, such as a random subset of the weight matrices, can yield strong performance while minimizing computational expenses.
 LoRA is simple, popular, and empirically effective for finetuning large language models.
-Many variants have been proposed to improve the initialization, performance, memory efficiency, or privacy of the algorithm [shen2025kroneckerlorahybridkroneckerloraadapters, büyükakyüz2024oloraorthonormallowrankadaptation, li2025flatlora, meng2024pissa, wang2024loraga], including but not limited to federated LoRA [10.24963/ijcai.2025/1196], quantized LoRA (QLoRA) [dettmersqlora], and ReLoRA [lialin2024relora].
-While LoRA performs well in practice, its behavior in comparison to full-rank training is poorly understood, with empirical studies showing that the two approaches can produce completely different solutions [shuttleworth2025loravsfinetuningillusion, biderman2024lora].
+Many variants have been proposed to improve the initialization, performance, memory efficiency, or privacy of the algorithm (Shen, 2025; büyükakyüz2024oloraorthonormallowrankadaptation; Li et al., 2025; Meng et al., 2024; Wang et al., 2024), including but not limited to federated LoRA (Yang et al., 2025), quantized LoRA (QLoRA) (Dettmers et al., 2023), and ReLoRA (Lialin et al., 2024).
+While LoRA performs well in practice, its behavior in comparison to full-rank training is poorly understood, with empirical studies showing that the two approaches can produce completely different solutions (Shuttleworth et al., 2025; Biderman et al., 2024).
 Consequently, the theoretic properties of LoRA are highly relevant for understanding the strengths and weakness of this method.
 As highlighted in the introduction, the theory of LoRA convergence has been tackled from many angles.
-First, many works examine the behavior of LoRA in infinite regimes, whether by characterizing the kinds of solutions LoRA converges to at infinity or analyzing its convergence in the neural tangent kernel (NTK) regime, which defines the linearized training dynamics of an infinitely wide neural network [pmlr-v202-malladi23alazy].
-For example, [kim2025lora] argues that LoRA converges to a global minimum with high probability, as long as such a low-rank minimizer exists.
-The work [zeng2024the] studies the expressive power of the low-rank solutions achieved by LoRA.
-The work [jang2024lora] shows that LoRA has no spurious local minima in the NTK regime.
-The LoRA+ [hayou2024loraefficientlowrank] algorithm, which updates the adapter matrices with differently scaled learning rates, is motivated by an NTK convergence analysis.
-Finally, the gradient flow dynamics of LoRA have been studied specifically for matrix factorization [xu2025understandingGF].
+First, many works examine the behavior of LoRA in infinite regimes, whether by characterizing the kinds of solutions LoRA converges to at infinity or analyzing its convergence in the neural tangent kernel (NTK) regime, which defines the linearized training dynamics of an infinitely wide neural network (Malladi et al., 2023).
+For example, (Kim et al., 2025) argues that LoRA converges to a low-rank global minimum with high probability, assuming that one exists.
+The work (Zeng and Lee, 2024) studies the expressive power of the low-rank solutions achieved by LoRA.
+The work (Jang et al., 2024) shows that LoRA has no spurious local minima in the NTK regime.
+The LoRA+ (Hayou et al., 2024) algorithm, which updates the adapter matrices with differently scaled learning rates, is motivated by an NTK convergence analysis.
+Finally, the gradient flow dynamics of LoRA have been studied specifically for matrix factorization (Xu et al., 2025).
 These works do not establish a convergence rate for discrete-time LoRA for general functions.
 Second, there are memory-efficient LoRA-like algorithms with concrete convergence analyses that do not apply to the original LoRA formulation.
-These include GaLore, an algorithm which projects the gradient into a low-rank subspace but maintains a full-rank update of the parameters [zhao2024galore], LDAdam [robert2025ldadam], which performs adaptive optimization steps within lower dimensional subspaces, and Randomized Subspace Optimization (RSO) [chen2025arso].
-The LoRA variant LoRA-One incorporates information from the full gradient at initialization [zhang2025loraone], and it is analyzed only for Gaussian data and mean-squared loss.
-Finally, the variants RAC-LoRA [malinovsky2024randomizedasymmetricchainlora] and Bernoulli-LoRA [sokolov2025bernoulliloratheoreticalframeworkrandomized] only train one adapter matrix while freezing the other as a fixed projection matrix, essentially preserving Lipschitz smoothness while failing to capture the complex training dynamics of actual LoRA implementations.
+These include GaLore, an algorithm which projects the gradient into a low-rank subspace but maintains a full-rank update of the parameters (Zhao et al., 2024), LDAdam (Robert et al., 2025), which performs adaptive optimization steps within lower dimensional subspaces, and Randomized Subspace Optimization (RSO) (Chen et al., 2025).
+The LoRA variant LoRA-One incorporates information from the full gradient at initialization (Zhang et al., 2025), and it is analyzed only for Gaussian data and mean-squared loss.
+Finally, the variants RAC-LoRA (Malinovsky et al., 2024) and Bernoulli-LoRA (Sokolov et al., 2025) only train one adapter matrix while freezing the other as a fixed projection matrix, essentially preserving Lipschitz smoothness while failing to capture the complex training dynamics of actual LoRA implementations.
 In practice, the LoRA matrices are updated simultaneously, causing nonlinearity and nonsmoothness in the optimized variables.
-As the original LoRA formulation is widely used in practice, these analyses do not address realistic implementations.
+As the original LoRA formulation is much more commonly used in practice, these analyses do not address realistic implementations.
 Third, several works prove the convergence of LoRA under a strong boundedness assumption that artificially enforces Lipschitz smoothness.
-This includes federated LoRA algorithms [sun2024improving, park2025communicationefficientfederatedlowrankupdate, ghiasvand2025decentralizedlowrankfinetuninglarge], for which single-device LoRA is a special case.
+This includes analyses of federated LoRA algorithms (Sun et al., 2024; Park and Klabjan, 2025; Ghiasvand et al., 2025), for which single-device LoRA is a special case.
 These works assume that the adapter matrix norms are uniformly bounded by constants.
-In addition, [jiang2024a] analyzes convergence of LoRA and representation fine-tuning (ReFT) within the same framework, assuming that the singular values of $A$ and $B$ are uniformly upper bounded, which upper bounds the matrix norms by rank-dependent constants that appear in the final convergence result.
+In addition, (Jiang et al., 2024) analyzes convergence of LoRA and representation fine-tuning (ReFT) within the same framework, assuming that the singular values of $A$ and $B$ are uniformly upper bounded, which upper bounds the matrix norms by rank-dependent constants that appear in the final convergence result.
 They also require that the gradient is uniformly bounded.
-These assumptions force Lipschitz smoothness in terms of the constants and are overly restrictive.
 Our work closes the gap in the existing literature by deriving the explicit convergence rate of the original LoRA algorithm for general models, without requiring bounded parameter norms or Lipschitz smoothness of the reparametrized loss function.
@@
 $\min_{W\in\mathbb{R}^{m\times n}}\ell(W_{0}+W),$ (1)
-where $\ell:\mathbb{R}^{m\times n}\to\mathbb{R}$ represents the loss function, $W_{0}\in\mathbb{R}^{m\times n}$ represents a frozen (pretrained) weight matrix, and $W\in\mathbb{R}^{m\times n}$ (sometimes denoted as $\Delta W$) represents the update to the frozen weights after fine-tuning.
+where $\ell:\mathbb{R}^{m\times n}\to\mathbb{R}$ represents the loss function, $W_{0}~\in~\mathbb{R}^{m\times n}$ represents a frozen (pretrained) weight matrix, and $W\in\mathbb{R}^{m\times n}$ (sometimes denoted as $\Delta W$) represents the update to the frozen weights after fine-tuning.
 In the LoRA algorithm, we parametrize $W=BA$, where $B\in\mathbb{R}^{m\times r}$, $A\in\mathbb{R}^{r\times n}$ are the low-rank adapter matrices with rank $r<\min\{m,n\}$.
@@
 While any kind of optimization can be applied to (2), in this work we consider the prevailing case of LoRA gradient descent, where at time step $t$ the matrices $A$, $B$ are updated simultaneously as follows,
-$A_{t+1}=A_{t}-\eta_{t}\nabla_{A}\mathcal{L}(B_{t}A_{t}),\qquad B_{t+1}=B_{t}-\eta_{t}\nabla_{B}\mathcal{L}(B_{t}A_{t}).$ (3)
-We require the following standard assumptions for our analysis.
-Assumption 1 (Lipschitz smoothness).
+$A_{t+1}=A_{t}-\eta_{t}\nabla_{A}\mathcal{L}(B_{t}A_{t}),$
+$B_{t+1}=B_{t}-\eta_{t}\nabla_{B}\mathcal{L}(B_{t}A_{t}),$ (3)
+where $\eta_{t}$ denotes the learning rate at time $t$. We require the following standard assumptions for our analysis.
+Assumption 3.1 (Lipschitz smoothness).
 The original loss function $\mathcal{L}:\mathbb{R}^{m\times n}\!\to\!\mathbb{R}$ is differentiable, and there exists a constant $L\geq 1$ such that for all $W,W^{\prime}\in\mathbb{R}^{m\times n}$,
 $\big\|\nabla\mathcal{L}(W)-\nabla\mathcal{L}(W^{\prime})\big\|_{F}\;\leq\;L\,\big\|W-W^{\prime}\big\|_{F}.$ (4)
-Equivalently, $\mathcal{L}$ satisfies the following “descent lemma”
+Equivalently, $\mathcal{L}$ satisfies the following descent lemma,
 $\mathcal{L}(W^{\prime})\leq\mathcal{L}(W)+\langle\nabla\mathcal{L}(W),W^{\prime}-W\rangle_{F}+\frac{L}{2}\lVert W-W^{\prime}\rVert_{F}^{2}.$ (5)
-Assumption 2.
+Assumption 3.2.
 The original loss function $\mathcal{L}$ is lower bounded by a constant $\mathcal{L}^{*}$ such that for all $W\in\mathbb{R}^{m\times n}$, $\mathcal{L}(W)\geq\mathcal{L}^{*}$.
-Assumptions 1 and 2 are fundamental to the study of gradient descent convergence.
+Assumptions 3.1 and 3.2 are fundamental to the study of gradient descent convergence.
 Critically, even with these assumptions, $\mathcal{L}$ is not Lipschitz smooth in $B$ or AA, preventing the analysis of (3) via classic optimization techniques.
@@
 We can rewrite the loss function in terms of the outer product $VV^{T}$.
-Second, we derive a modified “descent lemma,” analogous to the descent lemma for Lipschitz smooth functions (5), that enables descent in one step as long as the step size is small enough with respect to the parameter norm $\lVert V\rVert$ and gradient norm.
-The last step of the proof is ensuring that the step size does not decrease too quickly, thereby guaranteeing convergence.
+Second, we derive a modified descent lemma, analogous to the descent lemma for Lipschitz smooth functions (5), that enables descent in one step as long as the learning rate is small enough with respect to the parameter norm $\lVert V\rVert$ and gradient norm.
+The last step of the proof is ensuring that the learning rate does not decrease too quickly, thereby guaranteeing convergence.
 Step 1: Restructure problem into $VV^{T}$ form.
@@
 Let $I_{n}$ represent the $n\times n$ identity matrix.
-Let $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$ represent the “extractor” matrix that extracts the top $m$ rows from a $(m+n)\times(m+n)$ matrix when applied from the left, and let $E_{2}\penalty 10000\ =\penalty 10000\ \begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$ represent the matrix that extracts the right $n$ columns when applied from the right.
+Let $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$ represent the “extractor” matrix that extracts the top $m$ rows from a $(m+n)\times(m+n)$ matrix when applied from the left, and let $E_{2}~=~\begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$ represent the matrix that extracts the right $n$ columns when applied from the right.
 Then we have that
 $BA=E_{1}VV^{T}E_{2}.$
-We next define a new function $\mathcal{J}\penalty 10000\ :\penalty 10000\ \mathbb{R}^{(m+n)\times r}\to\mathbb{R}$ such that
+We next define the function $\mathcal{J}~:~\mathbb{R}^{(m+n)\times r}\to\mathbb{R}$ such that
 $\mathcal{J}(V)=\mathcal{L}(E_{1}VV^{T}E_{2})=\mathcal{L}(BA).$ (6)
@@
 $=$ $\begin{bmatrix}\nabla_{B}\mathcal{L}(BA)\\ \nabla_{A^{T}}\mathcal{L}(BA)\end{bmatrix}=\begin{bmatrix}\nabla_{B}\mathcal{L}(BA)\\ (\nabla_{A}\mathcal{L}(BA))^{T}\end{bmatrix}$.
-Step 2: “Lipschitz-like” descent lemma.
+Step 2: Descent lemma.
 We consider the gradient $\nabla\mathcal{J}(V)\in\mathbb{R}^{(m+n)\times r}$.
@@
 The function $\mathcal{J}$ is not Lipschitz smooth, due to the multiplicative factor of $V$ in $\nabla\mathcal{J}(V)$.
-However, we can still obtain a “descent lemma” (Lemma 1) that is analogous to the classic descent lemma of Lipschitz smooth functions (5), but with more higher-order terms.
-Lemma 1.
+However, we can still obtain a descent lemma (Lemma 3.3) that is analogous to the classic descent lemma of Lipschitz smooth functions (5), but with more higher-order terms.
+Lemma 3.3.
 For $\mathcal{J}(V)$ defined in (6), we have for all $V_{1},V_{2}\in\mathbb{R}^{(m+n)\times r}$,
-$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{\sqrt{2}L}{3}\lVert V_{2}-V_{1}\rVert^{3}$
-$+\frac{\sqrt{2}L}{4}\lVert V_{2}-V_{1}\rVert^{4}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
+$\mathcal{J}(V_{2})\leq$ $\mathcal{J}(V_{1})+\langle\nabla\mathcal{J}(V_{1}),V_{2}-V_{1}\rangle_{F}$
+$+\sqrt{2}L\lVert V_{2}-V_{1}\rVert^{2}\lVert V_{1}\rVert^{2}+\frac{\sqrt{2}L}{3}\lVert V_{2}-V_{1}\rVert^{3}$
+$+\frac{2\sqrt{2}}{3}L\lVert V_{2}-V_{1}\rVert^{3}\lVert V_{1}\rVert+\frac{\sqrt{2}L}{4}\lVert V_{2}-V_{1}\rVert^{4}$
+$+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
 Proof. See Appendix A.2.
-Step 3: Control step size to obtain convergence.
-Based on Lemma 1, we can pick $\eta_{t}$ small enough to minimize the higher-order terms, thereby guaranteeing descent in one step as stated next.
-Lemma 2.
+Step 3: Control learning rate to achieve convergence.
+Based on Lemma 3.3, we can pick $\eta_{t}$ small enough to minimize the higher-order terms, thereby guaranteeing descent in one step as stated next.
+Lemma 3.4.
 For any $V_{t}\in\mathbb{R}^{(m+n)\times r}$ and $V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, suppose that
@@
 Proof. See Appendix A.3.
-Lemma 2 states that to update $V_{t+1}$, we require the norm of the previous iterate $\lVert V_{t}\rVert$ and the gradient of the original loss function $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$.
+Lemma 3.4 states that to update $V_{t+1}$, we require the norm of the previous iterate $\lVert V_{t}\rVert$ and the gradient of the original loss function $\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$.
 By summing (9) over all t=0,,T1t=0,...,T-1 and telescoping sums, we determine that the minimum gradient norm over $t$ is upper bounded as
@@
 However, if the parameter norm is not bounded, then $\lVert V_{t}\rVert^{2}$ may grow to infinity.
-In the proof of Theorem 3, we show that even in the worst case, we have $\lVert V_{t}\rVert^{2}=O(t)$.
-The sum $\sum_{t=0}^{T-1}\eta_{t}$ is therefore lower bounded by a harmonic series which is $\Theta(\log(T))$, leading to Theorem 3.
-Theorem 3.
-Suppose Assumption 1 and 2 hold. We perform $T$ steps of LoRA gradient descent (3), where at time step $t$ the step size $\eta_{t}$ is determined as
+In the proof of Theorem 3.5, we show that even in the worst case, we have $\lVert V_{t}\rVert^{2}=O(t)$.
+The sum $\sum_{t=0}^{T-1}\eta_{t}$ is therefore lower bounded by a harmonic series which is $\Theta(\log(T))$, leading to Theorem 3.5.
+Theorem 3.5.
+Suppose Assumption 3.1 and 3.2 hold. We perform $T$ steps of LoRA gradient descent (3), where at time step $t$ the learning rate $\eta_{t}$ is determined as
 $\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\}.$
@@
 Proof. See Appendix A.4.
-Theorem 3 states that LoRA gradient descent converges to a stationary point at $O(\frac{1}{\log T})$ rate.
+Theorem 3.5 states that LoRA gradient descent converges to a stationary point at $O(\frac{1}{\log T})$ rate.
 If $\lVert V_{t}\rVert$ is uniformly bounded, the rate improves to $O(1/T)$.
@@
 However, even if $\lVert V_{t}\rVert$ grows and forces $\eta_{t}$ to decrease at each iteration, the algorithm still converges; a simple explanation is that if $\lVert V_{t}\rVert$ grows, $\eta_{t}$ decreases, which slows the growth of $\lVert V_{t}\rVert$.
-4 Discussion
+Extension to multiple matrices.
+The above theory assumes the loss functions is parametrized by a single weight matrix $W$.
+In practice, deep neural networks are parametrized by multiple weight matrices of potentially different sizes, each with its own LoRA adapter matrices.
+Our analysis naturally extends to this setting, which we illustrate (without loss of generality) with the following example of two matrices.
+Let $\mathcal{L}(W_{1},W_{2})\in\mathbb{R}$ denote the loss as a function of $W_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ and $W_{2}\in\mathbb{R}^{m_{2}\times n_{2}}$.
+For any matrix $M\in\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}$, define the linear operators $E_{11}(M)\in\mathbb{R}^{m_{1}\times n_{1}}$ and $E_{22}(M)\in\mathbb{R}^{m_{2}\times n_{2}}$ as follows,
+$E_{11}(M)=\begin{bmatrix}I_{m_{1}}&0_{m_{1}\times m_{2}}\end{bmatrix}M\begin{bmatrix}I_{n_{1}}\\ 0_{n_{2}\times n_{1}}\end{bmatrix}$,
+$E_{22}(M)=\begin{bmatrix}0_{m_{2}\times m_{1}}\times I_{m_{2}}\end{bmatrix}M\begin{bmatrix}0_{n_{1}\times n_{2}}\\ I_{n_{2}}\end{bmatrix}$,
+where $E_{11}$ extracts the top left $m_{1}\times n_{1}$ block and $E_{22}$ extracts the bottom right $m_{2}\times n_{2}$ block of $M$.
+We can define a new function $\tilde{\mathcal{L}}:\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}\to\mathbb{R}$ such that
+$\tilde{\mathcal{L}}(M)=\mathcal{L}(E_{11}(M),E_{22}(M)),$
+and the gradient $\nabla\tilde{\mathcal{L}}(M)$ is equal to the following matrix,
+$\begin{bmatrix}\nabla_{W_{1}}\mathcal{L}(E_{11}(M),E_{22}(M))&0\\ 0&\nabla_{W_{2}}\mathcal{L}(E_{11}(M),E_{22}(M))\end{bmatrix}$.
+This follows from the fact that by construction, the terms in the top right and bottom left blocks of $M$ have no impact on the value of $\tilde{\mathcal{L}}$, and their derivatives are zero.
+Overloading notation, we define the $\ell^{2}$ product norm on $\mathbb{R}^{m_{1}\times n_{1}}\times\mathbb{R}^{m_{2}\times n_{2}}$ as follows, for matrices $W_{1}\in\mathbb{R}^{m_{1}\times n_{1}}$ and $W_{2}\in\mathbb{R}^{m_{2}\times n_{2}}$,
+$\lVert(W_{1},W_{2})\rVert^{2}_{F}=\lVert W_{1}\rVert^{2}_{F}+\lVert W_{2}\rVert^{2}_{F}.$
+Then we can show that if $\mathcal{L}$ is Lipschitz smooth, $\tilde{\mathcal{L}}$ is also Lipschitz smooth.
+Lemma 3.6.
+Suppose $\mathcal{L}$ is $L$-Lipschitz smooth such that for all $W_{1},W_{1}^{\prime}\in\mathbb{R}^{m_{1}\times n_{1}}$, and $W_{2},W_{2}^{\prime}\in\mathbb{R}^{m_{2}\times n_{2}}$,
+$\lVert\nabla\mathcal{L}(W_{1},W_{2})-\nabla\mathcal{L}(W_{1}^{\prime},W_{2}^{\prime})\rVert\leq L\lVert(W_{1}-W_{1}^{\prime},W_{2}-W_{2}^{\prime})\rVert.$
+Then $\tilde{\mathcal{L}}$ is $2L$-Lipschitz smooth.
+Proof. See Appendix A.5.
+Finally, we reparametrize $W_{1}=B_{1}A_{1}$ and $W_{2}=B_{2}A_{2}$, where $B_{1}\in\mathbb{R}^{m_{1}\times r}$, $A_{1}\in\mathbb{R}^{r\times n_{1}}$, $B_{2}\in\mathbb{R}^{m_{2}\times r}$, $A_{2}\in\mathbb{R}^{r\times n_{1}}$ represent the LoRA adapter matrices.
+We can construct the larger low-rank matrices $B=[B_{1}^{T},B_{2}^{T}]^{T}$ and $A=[A_{1},A_{2}]$, such that
+$M=BA=\begin{bmatrix}B_{1}A_{1}&B_{1}A_{2}\\ B_{2}A_{1}&B_{2}A_{2}\end{bmatrix}$.
+Then $B_{1}A_{1}=E_{11}(BA)$ and $B_{2}A_{2}=E_{22}(BA)$, and gradient descent on $\tilde{L}(BA)$ with respect to $B$ and $A$ is equivalent to LoRA gradient descent on $\mathcal{L}$.
+Since $\tilde{L}$ is Lipschitz smooth, we can directly apply the existing analysis to this setting.
+Figure 1: Training logistic regression model on embeddings of the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt}$ and $\eta^{adapt2}$), and normalized ($\eta^{norm}$) learning rates.
+For clarity, the moving averages (window size $10$ and $200$ for a.
+and b.
+respectively) of the batch loss are plotted.
+3.3 Discussion
+Our theoretical results yield several interesting insights into the behavior of LoRA.
+First, the choice of extractor matrices $E_{1}$, $E_{2}$, has little to no impact on the proof, suggesting that other subsets of the $VV^{T}$ matrix could be extracted as alternate parameter-efficient reparametrizations to yield similar results.
+Moreover, the $VV^{T}$ form is exactly the symmetric Burer-Monteiro parametrization (Burer and Monteiro, 2003), and the results of our analysis can be trivially extended to show that gradient descent with Burer-Monteiro also converges to stationary points for general Lipschitz smooth functions.
+Second, the convergence rate does not explicitly depend on the chosen LoRA rank $r$, except that it is $O(\frac{1}{\lVert V_{0}\rVert^{2}})$.
+This is expected because $r$ just controls the dimension of the problem, but gradient descent is a dimension-free algorithm.
+Finally, due to the $\lVert V_{t}\rVert^{2}$ term in the denominator of $\eta_{t}$, the convergence of LoRA is slowed if the iterates are progressing away from the origin, and it is accelerated if they are moving towards the origin.
+This “position dependency” is not typical of gradient algorithms, and may be due to the fact that the LoRA reparametrization creates a stationary point at $V=0$ regardless of the structure of the original function.
+Consequently, LoRA may converge to the origin even if the original global minima is arbitrarily far away, explaining the observed divergence in behavior between LoRA and full-rank fine-tuning (Shuttleworth et al., 2025; Biderman et al., 2024).
+Figure 2: Training ResNet-18 model on the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt2}$) and normalized ($\eta^{norm}$) learning rates.
+For clarity, the moving averages (window size $50$) of the batch loss and gradient norm are plotted.
+4 Experiments
+We conduct experiments to investigate how the choice of learning rate and LoRA reparametrization affects training convergence in practice.
+We perform image classification on the CIFAR-10 dataset (Krizhevsky, 2009) with 10 classes, using the cross-entropy loss function.
+We first consider logistic regression, since the loss function is known to be Lipschitz smooth, and we train the model on embeddings of the CIFAR-10 images, produced by a ResNet-18 model (He et al., 2016) pretrained on ImageNet-1k (Deng et al., 2009).
+We then train a ResNet-18 model directly on the CIFAR-10 dataset, with LoRA implemented on the convolutional layers, and batch normalization layers disabled because they can have complex interactions with the learning rate.
+For each experimental setting, we replace the model weight matrices with low-rank approximations ($r=4$ and $r=20$) and we train with mini-batch SGD with a large batch size ($b=512$) to better capture our analysis of gradient descent.
+The model weights are first initialized randomly and frozen, and then the adapter matrices are initialized in the standard approach with $A$ as a random matrix, and $B$ as a zero matrix.
+For additional experimental details, including the anonymous GitHub repository, see Appendix B.
+We test three types of learning rate schemes: constant, adaptive, and normalized. We first consider an adaptive learning rate $\eta^{adapt}$ defined as follows,
+$\eta^{adapt}_{t}=\frac{\alpha}{\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}.$
+[…]
+$\eta_{t}^{adapt2}=\frac{\alpha}{\lVert V_{t}\rVert^{2}+\sqrt{\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})}}.$
+Finally, we also consider a normalized learning rate $\eta_{norm}$, defined as follows,
+$\eta^{norm}_{t}=\frac{\alpha}{\lVert\nabla\mathcal{L}(V_{t})\rVert^{1/2}}.$
+This inverse relationship between $\eta_{t}$ and $\lVert\nabla\mathcal{L}(V_{t})\rVert^{1/2}$ is required for minimizing higher order terms in the descent lemma, as stated in (21) and (23) in Appendix A.3.
+In practice, we compute $\eta^{adapt}_{t}$, $\eta^{adapt2}_{t}$ and $\eta^{norm}_{t}$ using the batch gradient and loss values at training step $t$.
+Figure 1 displays the results of the logistic regression experiments.
+Figures 1a and 1b demonstrate that training with the nonconstant learning rates $\eta_{t}^{adapt}$, $\eta_{t}^{adapt2}$, or $\eta_{t}^{norm}$ exhibits faster and more stable convergence than training with a constant learning rate in the same range of values.
+If the constant learning rate is too high, training becomes unstable, but if it is too low, convergence slows.
+Figure 1c displays the values of $\eta^{adapt}$, $\eta^{adapt2}$ and $\eta^{norm}$ over the training time, demonstrating that $\eta^{adapt}_{t}$ and $\eta^{adapt2}_{t}$ are closely correlated early on, but diverge as the model converges.
+Figure 2 displays the results of the ResNet-18 experiments using $\eta_{t}^{adapt2}$ and $\eta_{t}^{norm}$.
+Both the adaptive and normalized learning rates significantly stabilize training compared to constant learning rate while maintaining fast convergence, with $\eta^{norm}$ outperforming $\eta^{adapt2}$.
+Therefore, $\eta^{adapt2}$ and $\eta^{norm}$ are practical and computationally efficient methods for improving LoRA convergence on large neural networks.
+Figure 3: Training logistic regression model on embeddings of the CIFAR-10 dataset with constant, adaptive ($\eta^{adapt}$ and $\eta^{adapt2}$), and normalized ($\eta^{norm}$) learning rates.
+For clarity, the moving averages (window size $30$) of the batch gradient are plotted.
+Figures 3a and 2c display the progress of $\lVert V_{t}\rVert$ for various learning rates, which in combination with Figures 3b and 2d, suggests that the model is initialized in a flat region close to the origin, where the gradient norm is small.
+This aligns with the fact that LoRA always creates a stationary point at the origin.
+Our results suggest that a large initial learning rate is necessary for accelerating the iterates out of this region.
+Ultimately, our results demonstrate that $\eta^{adapt}$, $\eta^{adapt2}$, and $\eta^{norm}$ can improve LoRA training by taking advantage of the structural changes introduced by the low-rank reparametrization.
+These learning rate schemes start at higher values when the gradient norm is low and decrease over time, thereby accelerating training through the plateau around initialization and stabilizing it when the gradient norm is large.
+In particular, $\eta^{adapt}$ is most aligned with theory, while $\eta^{adapt2}$ and $\eta^{norm}$ are more appropriate for practical large-scale implementations.
+5 Conclusion
 In this work, we show for first time that the original LoRA gradient descent algorithm achieves a convergence rate of $O(\frac{1}{\log T})$, without requiring bounded adapter matrices or Lipschitz smoothness of the reparametrized loss function.
-Our results yield several interesting insights into the behavior of LoRA.
-First, the choice of extractor matrices $E_{1}$, $E_{2}$, has little to no impact on the proof, suggesting that other subsets of the $VV^{T}$ matrix could be extracted as alternate parameter-efficient reparametrizations to yield similar results.
-Moreover, the $VV^{T}$ form is exactly the symmetric Burer-Monteiro parametrization [burer_nonlinear_2003], and the results of our analysis can be trivially extended to show that gradient descent with Burer-Monteiro also converges to stationary points for general Lipschitz smooth functions.
-Second, the convergence rate does not explicitly depend on the chosen LoRA rank $r$, except that it is $O(\frac{1}{\lVert V_{0}\rVert^{2}})$.
-This is expected because $r$ just controls the dimension of the gradient descent algorithm, which is a dimension-free algorithm.
-Finally, our results shed light onto the complex relationship between the parameter norm and the LoRA training dynamics, which has been alluded to in past work.
-In particular, we observe that the convergence of LoRA is slowed if the iterates are progressing away from the origin, and it is accelerated if they are moving towards the origin.
-This “position dependency” is not typical of gradient algorithms, and may be due to the fact that the LoRA reparametrization creates a stationary point at $V=0$ regardless of the structure of the original function.
-Consequently, LoRA may converge to the origin even if the original global minima is arbitrarily far away, explaining the divergence in behavior between LoRA and full-rank fine-tuning.
 Future research directions may include determining convergence rates on convex or strongly convex loss functions, whose geometric properties may change under the LoRA reparametrization, as well as analyzing the stochastic gradient descent setting.
 A major challenge of this setting is deriving a fourth-moment bound on the noise.
+Impact Statement
+This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
 Appendix A Proofs
 A.1 Helper Lemmas
-Lemma 4.
+Lemma A.1.
 (Weighted) AM-GM inequality. Given the nonnegative values $x_{1},x_{2},w_{1},w_{2}$, we have
@@
 $x_{1}+x_{2}\geq 2\sqrt{x_{1}x_{2}}.$ (14)
-For the proof of Lemma 4, see Section 2.5 of [hardy_inequalities_1934].
-Lemma 5.
+For the proof of Lemma A.1, see Section 2.5 of (Hardy et al., 1934).
+Lemma A.2.
 For any matrix $A\in\mathbb{R}^{(m+n)\times(m+n)}$ and the extractor matrices $E_{1}=\begin{bmatrix}I_{m}&0_{m\times n}\end{bmatrix}$, $E_{1}\in\mathbb{R}^{m\times(m+n)}$, and $E_{2}=\begin{bmatrix}0_{n\times m}&I_{n}\end{bmatrix}^{T}$, $E_{2}\in\mathbb{R}^{(m+n)\times n}$
@@
 Then $\lVert A\rVert^{2}=\lVert A_{11}\rVert^{2}+\lVert A_{22}\rVert^{2}+2\lVert A_{12}\rVert^{2}$ and $\lVert A_{12}\rVert\leq\frac{1}{\sqrt{2}}\lVert A\rVert$. ∎
-Lemma 6.
+Lemma A.3.
 Let $\{a_{n}\}_{n\geq 0}$ be a real, nonnegative sequence, and let $c>0$ be a fixed constant.
@@
 $\sum_{n=1}^{\infty}a_{n}=\sum_{n=1}^{N-1}a_{n}+\sum_{n=N}^{\infty}a_{n}=\sum_{n=1}^{N-1}a_{n}+\sum_{n=N}^{\infty}b_{n}.$
-A.2 Proof of Lemma 1
+A.2 Proof of Lemma 3.3
 Proof.
@@
 $+\frac{\sqrt{2}}{4}L\lVert V_{2}-V_{1}\rVert^{4}+\lVert\nabla\mathcal{L}(E_{1}V_{1}V_{1}^{T}E_{2})\rVert\lVert V_{2}-V_{1}\rVert^{2}.$
-A.3 Proof of Lemma 2
+A.3 Proof of Lemma 3.4
 Proof.
-We apply the descent lemma (Lemma 1) with $V_{1}=V_{t}$, $V_{2}=V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, which yields
+We apply the descent lemma (Lemma 3.3) with $V_{1}=V_{t}$, $V_{2}=V_{t+1}=V_{t}-\eta_{t}\nabla\mathcal{J}(V_{t})$, which yields
 $\mathcal{J}(V_{t+1})\leq$ $\mathcal{J}(V_{t})-\eta_{t}\langle\nabla\mathcal{J}(V_{t}),\nabla\mathcal{J}(V_{t})\rangle_{F}+\frac{2\sqrt{2}}{3}L\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}\lVert V_{t}\rVert+\sqrt{2}L\eta_{t}^{2}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\lVert V_{t}\rVert^{2}+\frac{\sqrt{2}L}{3}\eta_{t}^{3}\lVert\nabla\mathcal{J}(V_{t})\rVert^{3}$
@@
 $\lVert\nabla\mathcal{J}(V)\rVert\leq 2\lVert\nabla\mathcal{L}(E_{1}VV^{T}E_{2})\rVert\lVert V\rVert.$ (19)
-Therefore, from (14) of Lemma 4, we have $\frac{1}{a+b}\leq\frac{1}{2\sqrt{ab}}$ for $a,b>0$, such that $\eta_{t}$ satisfies the following bound,
+Therefore, from (14) of Lemma A.1, we have $\frac{1}{a+b}\leq\frac{1}{2\sqrt{ab}}$ for $a,b>0$, such that $\eta_{t}$ satisfies the following bound,
 $\eta_{t}\leq\frac{1}{5\sqrt{2}L\lVert V_{t}\rVert^{2}+5\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}\leq$ $\left(\frac{1}{100\sqrt{2}L\lVert V_{t}\rVert^{2}\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert}\right)^{1/2},$
@@
 $\leq$ $\left(\frac{3}{10\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert\lVert V_{t}\rVert}\right)^{1/2}.$ (21)
-For the next bound, by (13) of Lemma 4 with $w_{1}=1$, $w_{2}=2$, $x_{1}=5\sqrt{2}\lVert V_{t}\rVert^{2}$, $x_{2}=\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$, we have
+For the next bound, by (13) of Lemma A.1 with $w_{1}=1$, $w_{2}=2$, $x_{1}=5\sqrt{2}\lVert V_{t}\rVert^{2}$, $x_{2}=\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert$, we have
 $5\sqrt{2}L\lVert V_{t}\rVert^{2}+2\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert\geq$ $3\left(5\sqrt{2}L\lVert V_{t}\rVert^{2}\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert^{2}\right)^{1/3},$
@@
     $\leq$ $\left(\frac{3}{5\sqrt{2}L\lVert\nabla\mathcal{J}(V_{t})\rVert}\right)^{1/2}.$
-In conclusion, with the choice of $\eta_{t}$ in (8), we can substitute the derived bounds on $\eta_{t}$ into Lemma 1, achieving descent in one step.
+In conclusion, with the choice of $\eta_{t}$ in (8), we can substitute the derived bounds on $\eta_{t}$ into Lemma 3.3, achieving descent in one step.
 […]
@@
 $\leq$ $\mathcal{J}(V_{t})-\frac{\eta_{t}}{5}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}.$
-A.4 Proof of Theorem 3
+A.4 Proof of Theorem 3.5
 Proof.
-From Lemma 2, we know that if
+From Lemma 3.4, we know that if
 $\eta_{t}=\min\{\frac{1}{5\sqrt{2}L(\lVert V_{t}\rVert^{2}+\lVert\nabla\mathcal{L}(E_{1}V_{t}V_{t}^{T}E_{2})\rVert)},1\},$
@@
 So to show convergence to a stationary point, we need to show that the sum of learning rates $\sum_{t=0}^{\infty}\eta_{t}$ diverges.
-From Lipschitz smoothness of $\mathcal{L}$, we have for all $W\in\mathbb{R}^{m\times n}$ (See Lemma 2.28 in [garrigos2024handbookconvergencetheoremsstochastic]),
-$\lVert\nabla\mathcal{L}(W)\rVert^{2}\leq$ $2L(\mathcal{L}(W)-\mathcal{L}^{*})$
-$\lVert\nabla\mathcal{L}(W)\rVert\leq$ $(2L(\mathcal{L}(W)-\mathcal{L}^{*}))^{1/2}$
+From Lipschitz smoothness of $\mathcal{L}$, we have for all $W\in\mathbb{R}^{m\times n}$ (See Lemma 2.28 in (Garrigos and Gower, 2024)),
+$\lVert\nabla\mathcal{L}(W)\rVert^{2}\leq 2L(\mathcal{L}(W)-\mathcal{L}^{*}),$ (24)
+which allows us to bound the gradient as follows,
+$\lVert\nabla\mathcal{L}(W)\rVert\leq$ $(2L(\mathcal{L}(W)-\mathcal{L}^{*}))^{1/2},$
 $\lVert\nabla\mathcal{L}(E_{1}VV^{T}E_{2})\rVert\leq$ $(2L(\mathcal{L}(E_{1}VV^{T}E_{2})-\mathcal{L}^{*}))^{1/2}=(2L(\mathcal{J}(V)-\mathcal{L}^{*}))^{1/2}.$
@@
 To show the divergence of $\sum_{t=0}^{\infty}\eta_{t}$, it is sufficient to show the divergence of the series $\sum_{t=0}^{\infty}\min\{\frac{1}{D_{t}},1\}$.
-Moreover, by the contrapositive of Lemma 6, we only need to show divergence of $\sum_{t=0}^{\infty}\frac{1}{D_{t}}$.
+Moreover, by the contrapositive of Lemma A.3, we only need to show divergence of $\sum_{t=0}^{\infty}\frac{1}{D_{t}}$.
 We first show that we achieve divergence if the adapter norms are bounded.
@@
 $\eta:=\min\{\frac{1}{5\sqrt{2}L(C^{2}+(2L(\mathcal{J}(V_{0})-\mathcal{L}^{*}))^{1/2})},1\}\leq\eta_{t},$
-and we prove (12) of Theorem 3 as follows
+and we prove (12) of Theorem 3.5 as follows
 $\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\eta T}=O(1/T).$
@@
 $\min_{t=0,...,T-1}\lVert\nabla\mathcal{J}(V_{t})\rVert^{2}\leq$ $\frac{5(\mathcal{J}(V_{0})-\mathcal{L}^{*})}{\sum_{t=0}^{T-1}\eta_{t}}=O\left(\frac{1}{\log T}\right).$
+A.5 Proof of Lemma 3.6
+Suppose $M,M^{\prime}\in\mathbb{R}^{(m_{1}+m_{2})\times(n_{1}+n_{2})}$. Then we have
+$\lVert\nabla\tilde{L}(M)-\nabla\tilde{L}(M^{\prime})\rVert^{2}=$ $\lVert\nabla_{W_{1}}\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla_{W_{1}}\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
+$+\lVert\nabla_{W_{2}}\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla_{W_{2}}\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
+$=$ $\lVert\nabla\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert^{2},$
+$\lVert\nabla\tilde{L}(M)-\nabla\tilde{L}(M^{\prime})\rVert=$ $\lVert\nabla\mathcal{L}(E_{11}(M),E_{22}(M))-\nabla\mathcal{L}(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert,$
+$\leq$ $L\lVert(E_{11}(M),E_{22}(M))-(E_{11}(M^{\prime}),E_{22}(M^{\prime}))\rVert,$
+$=$ $L\lVert(E_{11}(M-M^{\prime}),E_{22}(M-M^{\prime}))\rVert,$
+$\leq$ $L\lVert E_{11}(M-M^{\prime})\rVert+L\lVert E_{22}(M-M^{\prime})\rVert\leq 2L\lVert M-M^{\prime}\rVert,$
+where the last step follows from similar logic as Lemma A.2.
+Table 1: Experiment hyperparameters.
+Hyperparameter Logistic regression with ResNet-18 embeddings ResNet-18
+Batch size 512 512
+LoRA rank 4 20
+Data input dimension 512 $32\times 32\times 3$
+Epochs 60 250
+$\alpha^{adapt}$ 1
+$\alpha^{adapt2}$ 0.8 15
+$\alpha^{norm}$ 0.06 0.1
+Constant learning rate range [0.03, 0.05, 0.07] [0.03, 0.04, 0.045]
+Appendix B Experimental Details
+In the following, we provide additional details about our experimental setup.
+As mentioned in the paper, for the ResNet-18 experiments we turn off BatchNorm layers because they effectively change the learning rate.
+This is achieved by setting the model to evaluation mode with model.eval() before training.
+Table 1 displays the hyperparameters used in our experimental results.
+Code is open-sourced at the anonymous Github repository https://anonymous.4open.science/r/lora-C709/.
+Learning rate.
+While standard LoRA implementations tend to scale by both a scaling factor and $\frac{1}{r}$, for simplicity we do not scale by the rank and just set the learning rate as a whole.
+To determine the range of learning rates to test in our experiments, we use a cyclical learning rate finder as introduced in (Smith, 2017) and further described in (Ahmed, 2025).
+We start with a small learning rate and train the model for a single epoch, repeating this with exponentially increasing learning rates.
+We plot the loss over the learning rates and pick learning rates in the range where the loss decreases the fastest.
+The resulting plots are available in Figure 4.
+Figure 4: Learning rate selection using a cyclical finder for both experimental settings.
+Hardware and software.
+All experiments were run using PyTorch 2.5.0 and CUDA 12.1, on an Intel(R) Xeon(R) Silver 4208 CPU (2.10GHz) with an NVIDIA RTX A6000 GPU.